A solvable model of scattering by Helmholtz resonator

Boris Pavlov University of Auckland, New Zealand

The analytic perturbation procedure is divergent on the continuous spectrum near the threshold of creation of resonances. The analytic perturbation procedure is divergent on the continuous spectrum near the threshold of creation of resonances. The analytic perturbation procedure is divergent on the continuous spectrum near the threshold of creation of resonances. For scattering problem on the resonator with a small opening we suggest a modified procedure which is convergent and permits to observe the creation of the resonance and to construct a convergent analytical perturbation procedure with use of the solvable model as a jump-start.